OpenStudy (ashontae19):

easy question i will give a medal a 9th grader can answer them

OpenStudy (anonymous):

Whats the question?

OpenStudy (ashontae19):

@OpenSessame

OpenStudy (anonymous):

The numbers must add to get the center number and multiply to get the last number

OpenStudy (ashontae19):

ok i have the answers i just want to know if im right

OpenStudy (anonymous):

What did you get?

OpenStudy (ashontae19):

a

OpenStudy (ashontae19):

@OpenSessame

OpenStudy (anonymous):

A is correct

OpenStudy (ashontae19):

http://prntscr.com/7o0qwe

OpenStudy (ashontae19):

the second one

OpenStudy (ashontae19):

i got d

OpenStudy (anonymous):

D is correct

OpenStudy (ashontae19):

u sure?

OpenStudy (ashontae19):

@ottn656

OpenStudy (ashontae19):

ANYONE????

OpenStudy (anonymous):

I believe so

OpenStudy (ashontae19):

http://prntscr.com/7o0tyr

OpenStudy (ashontae19):

http://prntscr.com/7o0w84 the third one

OpenStudy (anonymous):

The answer is A

OpenStudy (ashontae19):

http://prntscr.com/7o0yf0

OpenStudy (ashontae19):

@OpenSessame

OpenStudy (anonymous):

I no good at geometry lol

OpenStudy (ashontae19):

http://prntscr.com/7o128w

OpenStudy (ashontae19):

what about this one

OpenStudy (ashontae19):

@DarkMoonZ

OpenStudy (ashontae19):

@welshfella

OpenStudy (welshfella):

the union of the 2 sets contains all the elements of the 2 sets ( duplicates not included)

OpenStudy (welshfella):

its not C because there is no 6

OpenStudy (ashontae19):

http://prntscr.com/7o1cfn what about this one

OpenStudy (welshfella):

the key word here is OR - OR means ADD the probabilities P(<4) = 2/5 P(even) = 1/5

OpenStudy (ashontae19):

it not 3/5?

OpenStudy (welshfella):

3/5 yes

OpenStudy (ashontae19):

http://prntscr.com/7o1ep5 what about this one?